# Mirror-Consistency: Harnessing Inconsistency in Majority Voting

### v1

@@ Abstract
Siyuan Huang1, [^1] Work done during an internship at Ant Group., Zhiyuan Ma2, Jintao Du2 Changhua Meng2, Weiqiang Wang2, Zhouhan Lin 1, [^2] Corresponding Author. 1 Shanghai Jiaotong University 2 Tiansuan Lab, Ant Group Co., Ltd. siyuan_huang_sjtu@outlook.com, lin.zhouhan@gmail.com {mazhiyuan.mzy,lingke.djt,changhua.mch,weiqiang.wwq}@antgroup.com
@@ 2 Mirror-Consistency
Mirror-Consistency and Self-Consistency share procedural similarities, as both methods generate multiple responses and use majority voting to determine the final output. The key distinction is that Mirror-Consistency examines the inconsistencies within the sampling process, rather than merely repeat sampling independent answers. The Mirror-Consistency process begins with an initial sampling phase to produce the first response, denoted as $r_{0}$. This is followed by a series of iterations that alternate between Reflection on Inconsistency and Conditional Resampling, as shown in Algorithm 1. We next introduce these two steps in detail. The corresponding prompts are provided in Appendix D.
@@
2:$r_{0}\leftarrow\mathcal{M}(p_{\text{sample}}\|x)$triangleright Initial response
3:$C_{0}\leftarrow\emptyset$triangleright Initial checklist
4:$m_{0}\leftarrow r_{0}$triangleright Initial majority vote
@@ Appendix A Additional Calibration Results
Here, $|\mathbf{\hat{r}}|$ represents the number of unique answers, and $p_{i}$ is the normalized frequency of each unique answer $\hat{r}_{i}$ in the dataset. This formulation inversely relates the entropy measure to the consistency of the answer distribution, where a lower entropy implies a more uniform and certain response pattern.
@@
[…]

### v2

@@ Abstract
Siyuan Huang1,3, [^1] Work done during an internship at Ant Group., Zhiyuan Ma2, Jintao Du2 Changhua Meng2, Weiqiang Wang2, Zhouhan Lin 1, [^2] Corresponding Author. 1 Shanghai Jiaotong University 2 Tiansuan Lab, Ant Group Co., Ltd. 3 SJTU Paris Elite Institute of Technology siyuan_huang_sjtu@outlook.com, lin.zhouhan@gmail.com {mazhiyuan.mzy,lingke.djt,changhua.mch,weiqiang.wwq}@antgroup.com
@@ 2 Mirror-Consistency
Mirror-Consistency and Self-Consistency share procedural similarities, as both methods generate multiple responses and use majority voting to determine the final output. The key distinction is that Mirror-Consistency examines the inconsistencies within the sampling process, rather than merely repeat sampling independent answers. The Mirror-Consistency process begins with an initial sampling phase to produce the first response, denoted as $r_{0}$. This is followed by a series of iterations that alternate between Reflection on Inconsistency and Conditional Resampling, as shown in Algorithm˜1. We next introduce these two steps in detail. The corresponding prompts are provided in Appendix D.
@@
[…]
@@ Appendix A Additional Calibration Results
[…]
@@
Again we derive a set of unique answers from the model’s output, denoted as $\mathbf{\hat{r}}$. The answer-number-based , $\texttt{Ans-Num}(\bar{r})$, is then simply defined as:
$\texttt{Ans-Num}(\bar{r})=1-\frac{|\mathbf{\hat{r}}|}{n}$